# Dynamics on Hyperspace of Pointwise Periodic Homeomorphisms

### v2

@@ Abstract.
In this paper, we consider the dynamics of induced map $2^{f}$ of a given pointwise periodic homeomorphism $f:X\to X$ of a compact metric space $X$. First, we show that the topological entropy of $2^{f}$ is zero, i.e. $h_{top}(2^{f})=0$ and that the set of almost periodic points coincides with the set of uniformly recurrent points, i.e. $AP(2^{f})=UR(2^{f})$. Furthermore, we prove that inside any infinite $\omega$-limit set $\omega_{2^{f}}(A)$ there is a unique minimal set and this minimal set is an adding machine. As a consequence, $(2^{X},2^{f})$ has no Devaney chaotic subsystems. In contrast to these rigidity properties, we obtain some results with chaotic flavor. In fact, we prove the following dichotomy, the hyperspace system $(2^{X},2^{f})$ is either equicontinuous or choatic with respect to Li-Yorke chaos and $\omega$-chaos. It is shown that the later case occurs if and only if $R(2^{f})\setminus AP(2^{f})\neq\emptyset$. This enables us to provide simple examples of pointwise periodic homeomorphisms with chaotic induced systems.
@@ 1. Introduction
This paper is organized as follow: In this section, we introduce the basic notations, definitions, and preliminary results that will be used in subsequent sections. In Section 2, we establish that the topological entropy $h_{top}(2^{f})$ is zero for any pointwise uniformly recurrent homeomorphism $f:X\to X$ of a compact metric space $X$. Section 3 is devoted to the study of pointwise periodic closed sets with the weak incompressibility property. In sections 4, 5 and 6, we will consider a pointwise periodic homeomorphism $f:X\to X$ of a compact metric space $X$. In section 4, we mainly prove that for an almost periodic point $A$ in the hyperspace $2^{X}$, $\cup_{n\in\mathbb{N}f^{n}(A)}$ is closed in $X$ and $A$ is uniformly recurrent for $2^{f}$. We prove in section 5, that inside any infinite $\omega$-limit set of $2^{f}$ there is a unique minimal set and this minimal set is an adding machine. Finally in section 6, we establish the following dichotomy result: $(2^{X},2^{f})$ is either equicontinuous or choatic with respect Li-Yorke chaos and $\omega$-chaos. The later case occurs if and only if $R(2^{f})\setminus AP(2^{f})\neq\emptyset$. We finish this section by giving a non-equicontinuous pointwise periodic homeomorphism $F:Y\to Y$ for which we construct a set $D\in 2^{Y}\setminus R(2^{F})$ and a set $C\in R(2^{F})\setminus AP(2^{F})$.
@@
  • transitive if $\omega_{f}(x)=X$.
@@ 2. Topological entropy of $2^{f}$
In this section we will prove that $h_{top}(2^{f})=0$ for any pointwise uniformly recurrent homeomorphism $f:X\to X$ of a compact metric space $X$. To achieve this, we will first show that the stable set with respect to $2^{f}_{|\omega_{2^{f}}(A)}$ of any recurrent point $A\in 2^{X}$ reduces to a single point set. Having shown this, we then invoke the variational principle together with the result of Blanchard, Host, and Ruelle (see [5]), which states that if the metric entropy relative to some ergodic probability measure $\mu$ is positive, then the set of points belonging to a proper asymptotic pair is of measure one with respect to $\mu$.
@@
Let $f:X\to X$ be a pointwise uniformly recurrent homeomorphism of a compact metric space $(X,d)$. Then $h_{top}(2^{f})=0$.
@@
Let $f:X\to X$ be a pointwise uniformly recurrent homeomorphism of a compact metric space $X$ and let $A,B\in 2^{X}$. If $A\in R(2^{f})$ and $(A,B)$ is an asymptotic pair of $2^{f}$ then $B\subset A$.
@@
Let $A,B\in 2^{X}$ be such that $A\in R(2^{f})$ and $(A,B)$ is an asymptotic pair of $2^{f}$. Suppose that $B\setminus A$ is not empty. Take $b\in B\setminus A$ and choose $\delta>0$ so that $B(b,2\delta)\subset X\setminus A$. As $b$ is uniformly recurrent for $f$, there exists $N\in\mathbb{N}$ such that $d(b,f^{kN}(b))<\delta$, $\forall k\in\mathbb{N}$. Now as $A\in R(2^{f^{N}})$, then $d_{H}(A,f^{kN}(A))<\delta/2$ for infinitely many $k\in\mathbb{N}$. Note further that for $k$ large enough, $d_{H}(f^{kN}(A),f^{kN}(B))<\delta/2$. It turns out that $d_{H}(f^{kN}(B),A)<\delta$ for infinitely many $k$. Consequently, for infinitely many $k\in\mathbb{N}$
$d(b,A)\leq d(b,f^{kN}(b))+d(f^{kN}(b),A)<\delta+d_{H}(f^{kN}(B),A)<2\delta.$
This is a contradiction with the fact that $d(b,A)\geq 2\delta$. Thus $B\setminus A=\emptyset$ and so $B\subset A$. ∎
@@
Let $f:X\to X$ be a pointwise uniformly recurrent homeomorphism of a compact metric space $X$. Then for any $A\in R(2^{f})$,
@@
Let $f:X\to X$ be a pointwise uniformly recurrent homeomorphism of a compact metric space $X$. Take $\mu$ an ergodic measure of $2^{f}$ and denote by $S$ its support. Clearly if $S$ is finite then the metric entropy $h_{\mu}(2^{f})=0$. In view of Poincaré’s Recurrence Theorem, any isolated point $Z$ of $S$ is periodic and so it has finite orbit with positive $\mu$-measure. As $\mu$ is ergodic, its support $S$ should be reduced to the orbit of $Z$ if $S$ contains an isolated point $Z$ and so it is finite in this case. Therefore, if $S$ is infinite then it has no isolated point. Hence the set of transitive points for the dynamical system $(S,2^{f}_{\mid S})$ has full measure with respect to $\mu$ (see Theorem 5.16 in [22]). Thus for $\mu$-a.e. $A\in 2^{X}$, $\omega_{2^{f}}(A)=S$ and $A\in S$. Consequently, for $\mu$-a.e. $A\in 2^{X}$, $A\in R(2^{f})$ and $\omega_{2^{f}}(A)=S$ thus $W^{s}(A)\cap S=\{A\}$ by Proposition 2.3. According to Blanchard, Host and Ruette’s result in [5], $h_{\mu}(2^{f})=h_{\mu}(2^{f}_{|S})=0$. Finally the topological entropy $h_{top}(2^{f})$ vanishes by the variational principle Theorem (see Corollary 8.6.1 in [22]). ∎
@@
In a joint work [8] with A. Daghar, we provided an example of a pointwise periodic rational curve homeomorphism $F$ and we claimed that $h_{top}(2^{F})=+\infty$(see Example 4.9 in [8]). It is worth noticing that the entropy of $2^{F}$ was incorrectly computed. In fact according to the above result, it vanishes. However, as we will see later in Section 5, the induced map $2^{F}$ of this particular example still exhibits chaotic features, such as Li-Yorke chaos.

### v3

@@ Abstract.
In this paper, we first prove that the topological entropy of induced map of any distal homeomorphism of a compact metric space is null. Then we consider induced map $2^{f}$ of an arbitrary pointwise periodic homeomorphism $f:X\to X$ of a compact metric space $X$, we show that the set of almost periodic points coincides with the set of uniformly recurrent points, i.e. $AP(2^{f})=UR(2^{f})$. Furthermore, we prove that inside any infinite $\omega$-limit set $\omega_{2^{f}}(A)$ there is a unique minimal set and this minimal set is an adding machine. As a consequence, $(2^{X},2^{f})$ has no Devaney chaotic subsystems. In contrast to these rigidity properties, we obtain some results with chaotic flavor. In fact, we prove the following dichotomy, the hyperspace system $(2^{X},2^{f})$ is either equicontinuous or choatic with respect to Li-Yorke chaos and $\omega$-chaos. It is shown that the later case occurs if and only if $R(2^{f})\setminus AP(2^{f})\neq\emptyset$. This enables us to provide simple examples of pointwise periodic homeomorphisms with chaotic induced systems.
@@ 1. Introduction
This paper is organized as follow: In this section, we introduce the basic notations, definitions, and preliminary results that will be used in subsequent sections. In Section 2, we establish that the topological entropy $h_{top}(2^{f})$ is zero for any distal homeomorphism $f:X\to X$ of a compact metric space $X$. Section 3 is devoted to the study of pointwise periodic closed sets with the weak incompressibility property. In sections 4, 5 and 6, we will consider a pointwise periodic homeomorphism $f:X\to X$ of a compact metric space $X$. In section 4, we mainly prove that for an almost periodic point $A$ in the hyperspace $2^{X}$, $\cup_{n\in\mathbb{N}f^{n}(A)}$ is closed in $X$ and $A$ is uniformly recurrent for $2^{f}$. We prove in section 5, that inside any infinite $\omega$-limit set of $2^{f}$ there is a unique minimal set and this minimal set is an adding machine. Finally in section 6, we establish the following dichotomy result: $(2^{X},2^{f})$ is either equicontinuous or choatic with respect Li-Yorke chaos and $\omega$-chaos. The later case occurs if and only if $R(2^{f})\setminus AP(2^{f})\neq\emptyset$. We finish this section by giving a non-equicontinuous pointwise periodic homeomorphism $F:Y\to Y$ for which we construct a set $D\in 2^{Y}\setminus R(2^{F})$ and a set $C\in R(2^{F})\setminus AP(2^{F})$.
@@
  • transitive if $\omega_{f}(x)=X$;
  • distal if $(x,y)$ is a distal pair for any $y\in\overline{O_{f}(x)}\setminus\{x\}$.
@@ 2. Topological entropy of $2^{f}$
In this section we will prove that $h_{top}(2^{f})=0$ for any distal homeomorphism $f:X\to X$ of a compact metric space $X$. To achieve this, we will first show that the stable set with respect to $2^{f}_{|\omega_{2^{f}}(A)}$ of any recurrent point $A\in 2^{X}$ reduces to a single point set.This was done using the product recurrence property of distal points: If $x$ is a distal point in a dynamical system $(X,T)$ then for any dynamical system $(Y,S)$ and any recurrent point $y\in Y$, the pair $(x,y)$ is a recurrent point in the product system $(X\times Y,T\times S)$ (see [9, Theorem 9.11]). Having shown this, we then invoke the variational principle together with the result of Blanchard, Host, and Ruelle (see [5]), which states that if the metric entropy relative to some ergodic probability measure $\mu$ is positive, then the set of points belonging to a proper asymptotic pair is of measure one with respect to $\mu$.
@@
Let $f:X\to X$ be a distal homeomorphism of a compact metric space $(X,d)$. Then $h_{top}(2^{f})=0$.
@@
Let $f:X\to X$ be a distal homeomorphism of a compact metric space $X$ and let $A,B\in 2^{X}$. If $A\in R(2^{f})$ and $(A,B)$ is an asymptotic pair of $2^{f}$ then $B\subset A$.
@@
Let $A,B\in 2^{X}$ be such that $A\in R(2^{f})$ and $(A,B)$ is an asymptotic pair of $2^{f}$. Suppose that $B\setminus A$ is not empty. Take $b\in B\setminus A$ and choose $\delta>0$ so that $B(b,2\delta)\subset X\setminus A$. As $b$ is a distal point, the pair $(b,A)$ is recurrent under the action of the homeomorphism $(f,2^{f})$. Thus for a suitable infinite sequence of positive integers $(n_{i})_{i}$, the limit in the product space $X\times 2^{X}$,
$\lim_{i\to+\infty}(f^{n_{i}}(b),f^{n_{i}}(A))=(b,A).$
However as $(A,B)$ is an asymptotic pair for $2^{f}$, we have also
$\lim_{i\to+\infty}(f^{n_{i}}(b),f^{n_{i}}(B))=(b,A).$
It follows that for $i$ large enough, $f^{n_{i}}(b)\in B(b,\delta)\cap B(A,\delta)$ which is clearly a contradiction. Consequently, $B\setminus A=\emptyset$ and so $B\subset A$. ∎
@@
Let $f:X\to X$ be a distal homeomorphism of a compact metric space $X$. Then for any $A\in R(2^{f})$,
@@
Let $f:X\to X$ be a distal homeomorphism of a compact metric space $X$. Take $\mu$ an ergodic measure of $2^{f}$ and denote by $S$ its support. Clearly if $S$ is finite then the metric entropy $h_{\mu}(2^{f})=0$. In view of Poincaré’s Recurrence Theorem, any isolated point $Z$ of $S$ is periodic and so it has finite orbit with positive $\mu$-measure. As $\mu$ is ergodic, its support $S$ should be reduced to the orbit of $Z$ if $S$ contains an isolated point $Z$ and so it is finite in this case. Therefore, if $S$ is infinite then it has no isolated point. Hence the set of transitive points for the dynamical system $(S,2^{f}_{\mid S})$ has full measure with respect to $\mu$ (see Theorem 5.16 in [24]). Thus for $\mu$-a.e. $A\in 2^{X}$, $\omega_{2^{f}}(A)=S$ and $A\in S$. Consequently, for $\mu$-a.e. $A\in 2^{X}$, $A\in R(2^{f})$ and $\omega_{2^{f}}(A)=S$ thus $W^{s}(A)\cap S=\{A\}$ by Proposition 2.3. According to Blanchard, Host and Ruette’s result in [5], $h_{\mu}(2^{f})=h_{\mu}(2^{f}_{|S})=0$. Finally the topological entropy $h_{top}(2^{f})$ vanishes by the variational principle Theorem (see Corollary 8.6.1 in [24]). ∎
@@
  • (i)
    The assumption of pointwise distality in Theorem 2.1 is essential. Indeed, there exist well-known examples of minimal Toeplitz subshifts with positive entropy that admit a generic set of distal points. Moreover, this result cannot be strengthened by passing to sequence topological entropy. Since as shown by Qiu and Zhao in [21], non-equicontinuous distal systems must have infinite sequence topological entropy.
  • (ii)
    In [8], we provided an example of a pointwise periodic rational curve homeomorphism $F$ and we claimed that $h_{top}(2^{F})=+\infty$ (see Example 4.9 in [8]). It is worth noticing that the entropy of $2^{F}$ was incorrectly computed. In fact according to the above result, it vanishes. However, as we will see later in Section 5, the induced map $2^{F}$ of this particular example still exhibits chaotic features, such as Li-Yorke chaos.